OpenStudy (anonymous):

simplify (csc⁡(θ) (1-〖cos⁡(θ)〗^2 ))/cos⁡(θ)

zepdrix (zepdrix):

Hey Timmy! Welcome to OpenStudy! \[\Large\rm \frac{\csc \theta(1-\cos^2\theta)}{\cos \theta}\]Is that what the expression looks like? Square on only the cosine?